10/34 . 17/19 . 19/20=


thats 10/34 * 17/19 * 19/20 =
can anyone help me out i mean i could get the answer but the numbers are way to large for me to factor step by step instructions as well as answer

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

10/34 * 17/19 * 19/20 =
= (10*17*19)/(34*19*20)
Cancel the 19s
= (10*17)/(34*20)
The 10s are obvious, cancel the 10 and 20 = 2*10
= 17/(34*2)
Cancel the 17s
= 1/4
Now don't you feel silly, talking about "large" numbers?

Answer by stanbon(75887)

10/34 * 17/19 * 19/20 =
---
10/(2*17) * (17/19) * 19/(2*10)
---
Cancel the 10's
Cancel the 17's
Cancel the 19's
-------
= 1/2 * 1/1 * 1/2
---
= 1/4
==============
Cheers,
Stan H.